OpenStudy (timaashorty):

Can anyone please help me ? Find a quadratic equation with roots -1+4i and -1-4i.

OpenStudy (anonymous):

there are a couple ways to do this, one is to work backwards \[x=-1+4i\] \[x+1=4i\] \[(x+1)^2=-16\] \[x^2+2x+1=-16\] \[x^2+2x+17=0\]

OpenStudy (anonymous):

another way is to memorize that if \(a+bi\) is the zero of a quadratic, then the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

in your case \(a=-1,b=4\) so you get \[x^2-2\times (-1)x+(1^2+4^2)=x^2-2x+17\] as before

OpenStudy (anonymous):

oops i meant \(x^2+2x+17\) as before

OpenStudy (timaashorty):

Oh my Thank you for explaing it step by step ! The final answer would be x^2+2x+17 right ?

OpenStudy (timaashorty):

@satellite73 ?